OpenStudy (anonymous):

Given right triangle ABC, where side "c" is the hypotenuse, angle A measures 40 degrees, and side b measures 17, find the length of side a.

OpenStudy (anonymous):

i know i closed the question, but i still need help with an answer, i have to get an A on this assignment.

OpenStudy (ranga):



OpenStudy (ranga):

tan(x) = opposite / adjacent tan(40) = BC / AC = a / 17 tan(40) = a/17 multiply both sides by 17 17 * tan(40) = a set your calculator to degrees. find tan(40) and multiply by 17

OpenStudy (anonymous):

14.26

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

can you help me with a few more, i need to get an a. and am running outta time. im only using this website for this quiz, the stuff after this material i have figured out. its this stuff im stuck on. youll never see me back again.